Ian 410109cle055d6pittmath1%2fdefault 127308 O E M
Answer #1

initial horizontal component of velocity Vx = V cos 45 = 510 * cos 45 = 360.57 m/s

initial vertical component of velocity Vy = V cos 45 = 510 * sin 45 = 360.57 m/s

use for vertical motion Y = Vyt -0.5 gt^2

4.9t^2 -360 .57 t -6 = 0

solving for t , t = 73.60 secs

now for horizontal motion, distance S = Vxt = 360.57 * 73.06

Sx = 26.53 km -----------<<<<<<<<<<<<Answer to part A

-------------------------------------------

vxf = u -gt = (360.57 - 721.23) = - 360.73 m/s

Vf^2 = Vx^2 + Vy^2

Vf^2 = 360.73^2 + (-360.73^2)

Vf = 510 m/s --------------------------<<<<<<<<<<<Answer to part B